A boy travels 10m due north and then 10m due east. What is the displacement of body?


1
Expert's answer
2021-09-15T17:32:31-0400

We are given that first a boy travels 10m

towards the north. We can take the north direction as the +y axis and so we can take this displacement in the north direction as 10jˆm

.After that he travels 7m

towards the east. We can take the east direction as the +x axis and so we can take this displacement in the east direction as 7iˆm

The total displacement is (7iˆ+10jˆ)m. We know that the magnitude of a vector quantity (xiˆ+yjˆ)is given by:

"\sqrt{\smash[b]{x^2 + y^2}}"

magnitude of the required displacement is given by:

= "\sqrt{\smash[b]{7^2+10^2}}" = "\sqrt{\smash[b]{49+100}}"

"= \sqrt{\smash[b]{149}}"

= 12.2m